HETAV PATEL

Grade 11,

2. Find the LCM and HCF of the following pairs of integers and verify that LCM × HCF = product of the two numbers. 510 and 92

2. Find the LCM and HCF of the following pairs of integers and verify that LCM × HCF = product of the two numbers.   510 and 92 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

expressing 510 and 92 as product of prime factors.

510 = 2*3*17*5*1

92 = 2*2*23*1

LCM= 23460
HCF= 2

Now, product of 510 and 92 = 510 *92= 46920
product of LCM and HCF = 23460*2= 46920
Hence, Proved.

Thanks